# Activation Oracles: Training and Evaluating LLMs as General-Purpose Activation Explainers

### v1

@@ Abstract
Large language model (LLM) activations are notoriously difficult to understand, with most existing techniques using complex, specialized methods for interpreting them. Recent work has proposed a simpler approach known as LatentQA: training LLMs to directly accept LLM activations as inputs and answer arbitrary questions about them in natural language. However, prior work has focused on narrow task settings for both training and evaluation. In this paper, we instead take a generalist perspective. We evaluate LatentQA-trained models, which we call Activation Oracles (AOs), in far out-of-distribution settings and examine how performance scales with training data diversity. We find that AOs can recover information fine-tuned into a model (e.g., biographical knowledge or malign propensities) that does not appear in the input text, despite never being trained with activations from a fine-tuned model. Our main evaluations are four downstream tasks where we can compare to prior white- and black-box techniques. We find that even narrowly-trained LatentQA models can generalize well, and that adding additional training datasets (such as classification tasks and a self-supervised context prediction task) yields consistent further improvements. Overall, our best AOs match or exceed prior white-box baselines on all four tasks and are the best method on 3 out of 4. These results suggest that diversified training to answer natural-language queries imparts a general capability to verbalize information about LLM activations.
@@ 1 Introduction
We find that narrowly-trained AOs can already succeed at these tasks and that scaling the quantity and diversity of training data yields consistent additional gains. We replicate this finding on three open-weight models (Qwen3-8B, Gemma-2-9B-IT, and Llama-3.3-70B-Instruct) and one closed-weight model (Claude Haiku 3.5). Overall, our best AOs match or exceed white-box baselines on all four downstream tasks and are the best overall method (including white- and black-box techniques) on 3 out of 4. This is especially striking because, once trained, AOs can be applied to these tasks out-of-the-box, without the task-specific scaffolding and tuning many other methods require.
@@
    We compare AOs to prior methods on downstream tasks, finding that they are the best-performing on 3 out of 4 of the tasks we study. They also match or exceed the performance of the best prior white-box techniques on all tasks. While prior techniques require task-specific scaffolding and tuning, AOs can be applied directly by extracting activations and asking natural-language questions.
@@ 3.1 Activation Steering for Latent Verbalization
Injection Mechanism. At layer 1 (see Appendix A.5 for layer selection details) of the forward pass, we modify the residual stream activations at each placeholder position by adding a norm-matched steering vector. For each placeholder token at position $i$ with corresponding injected vector $\mathbf{v}_{i}$:
@@ 3.3 Implementation Details
Training Scale. Our training dataset consists of approximately 1 million examples across all tasks. We train our Activation Oracles using LoRA adapters. The process is computationally inexpensive, requiring 10 H100 GPU hours for Qwen3-8B and 90 H200 hours for Llama-3.3-70B. For further training details and hyperparameters, refer to Appendix A.
@@ 4 Evaluation on downstream Auditing Tasks
Figure 3: Our Activation Oracle outperforms all existing white box baselines across all secret-keeping benchmarks. In these setups we take activations from the full sequence of tokens in the target prompt. The oracle prompt shows the natural language instructions that we provide to the AO. Despite training on very different tasks, the Activation Oracle generalizes and follows instructions to retrieve hidden behavior from these models.
@@ 4.1 Secret-Keeping Evaluations
Evaluation Setup. For the secret-keeping benchmarks, we use a simple approach: we collect activations from prompts that do not reference the secret information, then pose a single direct question to the Activation Oracle and check if the answer is correct. The target prompts are those used in cywiński2025elicitingsecretknowledgelanguage, which directly attempt to elicit the secret information; the models give benign responses to these prompts that do not reveal the secrets. For all secret-keeping evaluations, we use the full sequence of activations as input for simplicity and consistency, though we find that single-token inputs can improve performance on the Taboo task (see Appendix C.4).
We compare against existing baselines from cywiński2025elicitingsecretknowledgelanguage, which employ agent-based frameworks where auditing agents have access to five white-box interpretability tools (including logit lens, SAEs, and activation steering) and five black-box tools (including prompt injection and fuzzing). These baselines use manually-optimized pipelines for each task, which use from one to four steps.
@@
The performance gains vary across tasks. In some cases, the improvement is modest, such as achieving 59% versus 52% accuracy on the Secret Side Constraint task. In other cases, the gains are substantial: we achieve 21% accuracy on Taboo secret-keeping compared to just 4% for the best interpretability baseline. Additionally, Activation Oracles outperform the best black-box methods on two of the three tasks. These results are obtained using a considerably simpler method: we provide the Activation Oracle with a single prompt rather than employing the multi-stage pipelines developed by cywiński2025elicitingsecretknowledgelanguage.
@@ 10 Acknowledgements
We would like to thank Dan Mossing, Abhay Sheshadri, Jack Lindsey, and Daniel Filan for useful discussions and helpful feedback. We thank Andy Arditi for creating the PersonaQA dataset. AK did this work as part of the MATS Fellowship. JC and OE were supported by grants from Coefficient Giving. We are grateful to Constellation for hosting.
@@ A.2 Group by Length Batching
As we did not implement sequence packing, we instead use group by length batching to minimize padding tokens and improve training efficiency. We group training examples into mega-batches of size $\text{batch\_size}\times\text{window\_size}$ (16 x 20 = 320 examples with our settings), sort by sequence length in descending order within each mega-batch, then flatten back into the full training order. This ensures that examples within each training batch have similar lengths while maintaining sufficient randomness across mega-batches. We found this approach yielded a speedup of approximately 30% with no change to final train loss or evaluation performance.
@@ A.3 Infrastructure
We generate all activation vectors on the fly during training by temporarily disabling the LoRA adapter when running the target model forward pass. This avoids prohibitive storage costs for full sequences of activations.
@@ A.5 Steering Layer Selection
We investigated whether to apply activation steering at layer 0 (immediately after the embedding layer) or layer 1 of the Activation Oracle. While pan2024latentqateachingllmsdecode found layer 0 optimal, we hypothesized that layer 1 might work better with LoRA fine-tuning. The transition from layer 0 to layer 1 represents a large transformation in the model’s representation space, and we hypothesized that steering after this initial processing might be easier for the LoRA adapter to handle.
@@
Activation Norm Growth During Training. When using the replacement-based steering method from LatentQA, we observed problematic norm growth:
@@ C.4 Evaluation Input Token Selection
This suggests that providing the full sequence is a reasonable default, though having the flexibility to use single-token inputs can be valuable for specific tasks. However, in real-world auditing applications without ground-truth labels, it may be difficult to sweep token selection as a hyperparameter. In such cases, providing the full sequence of activations is the simplest option and performs well across most settings.

### v2

@@ Abstract
Large language model (LLM) activations are notoriously difficult to understand, with most existing techniques using complex, specialized methods for interpreting them. Recent work has proposed a simpler approach known as LatentQA: training LLMs to directly accept LLM activations as inputs and answer arbitrary questions about them in natural language. However, prior work has focused on narrow task settings for both training and evaluation. In this paper, we instead take a generalist perspective. We evaluate LatentQA-trained models, which we call Activation Oracles (AOs), in far out-of-distribution settings and examine how performance scales with training data diversity. We find that AOs can recover information fine-tuned into a model (e.g., biographical knowledge or malign propensities) that does not appear in the input text, despite never being trained with activations from a fine-tuned model. Our main evaluations are four downstream tasks where we can compare to prior white- and black-box techniques. We find that even narrowly-trained LatentQA models can generalize well, and that adding additional training datasets (such as classification tasks and a self-supervised context prediction task) yields consistent further improvements. Our best AOs match or exceed white-box baselines on all four tasks and the best overall baseline on 3 of 4. These results suggest that diversified training to answer natural-language queries imparts a general capability to verbalize information about LLM activations.
@@ 1 Introduction
We find that narrowly-trained AOs can already succeed at these tasks and that scaling the quantity and diversity of training data yields consistent additional gains. We replicate this finding on three open-weight models (Qwen3-8B, Gemma-2-9B-IT, and Llama-3.3-70B-Instruct) and one closed-weight model (Claude Haiku 3.5). Overall, our best AOs match or exceed white-box baselines on all four downstream tasks and match or exceed the best baseline (including white- and black-box techniques) on 3 out of 4. This is especially striking because, once trained, AOs can be applied to these tasks out-of-the-box, without the task-specific scaffolding and tuning many other methods require.
@@
    We compare AOs to prior methods on downstream tasks, finding that they match or exceed the best baseline method on 3 out of 4 of the tasks we study. They also match or exceed the performance of the best prior white-box techniques on all tasks. While prior techniques require task-specific scaffolding and tuning, AOs can be applied directly by extracting activations and asking natural-language questions.
@@ 3.1 Activation Steering for Latent Verbalization
Injection Mechanism. After the second transformer layer (see Appendix A.5 for layer selection details), we modify the residual stream activations at each placeholder position by adding a norm-matched steering vector. For each placeholder token at position $i$ with corresponding injected vector $\mathbf{v}_{i}$:
@@ 3.3 Implementation Details
Training Scale. Our training dataset consists of approximately 1 million examples (about 65M tokens) across all tasks. We train our Activation Oracles using LoRA adapters. The process is computationally inexpensive, requiring 10 H100 GPU hours for Qwen3-8B and 90 H200 hours for Llama-3.3-70B. For further training details and hyperparameters, refer to Appendix A.
@@ 4 Evaluation on downstream Auditing Tasks
Figure 3: Our Activation Oracle outperforms all existing white box baselines across all secret-keeping benchmarks. In these setups we take activations from the full sequence of tokens in the target prompt. The oracle prompt shows the natural language instructions that we provide to the AO. Despite training on very different tasks, the Activation Oracle generalizes and follows instructions to retrieve hidden behavior from these models. We use the baseline results from cywiński2025elicitingsecretknowledgelanguage (Appendix C.1).
@@ 4.1 Secret-Keeping Evaluations
Evaluation Setup. For the secret-keeping benchmarks, we use a simple approach: we collect activations from prompts that do not reference the secret information, then pose a single direct question to the Activation Oracle and check if the answer is correct. The target prompts are those used in cywiński2025elicitingsecretknowledgelanguage, which directly attempt to elicit the secret information; the models give benign responses to these prompts that do not reveal the secrets. For all secret-keeping evaluations, we use the full sequence of activations as input for simplicity and consistency, though we find that alternative token selections can improve performance, such as single-token inputs on the Taboo task and shorter selected sequences for Gender (see Appendix C.4).
We compare against existing baselines reported by cywiński2025elicitingsecretknowledgelanguage (Appendix C.1), which employ agent-based frameworks where auditing agents have access to five white-box interpretability tools (including logit lens, SAEs, and activation steering) and five black-box tools (including prompt injection and fuzzing). These baselines use manually-optimized pipelines for each task, which use from one to four steps.
@@
The performance gains vary across tasks. In some cases, the improvement is modest, such as achieving 59% versus 52% accuracy on the Secret Side Constraint task. In other cases, the gains are substantial: we achieve 21% accuracy on Taboo secret-keeping compared to just 4% for the best interpretability baseline. Additionally, Activation Oracles match or exceed the best black-box methods on two of the three tasks. These results are obtained using a considerably simpler method: we provide the Activation Oracle with a single prompt rather than employing the multi-stage pipelines developed by cywiński2025elicitingsecretknowledgelanguage.
@@ 9 Conclusion
10 Author Contributions
Adam Karvonen led the project, created the training infrastructure, implemented all experiments not otherwise mentioned, and wrote the initial draft of the paper with contributions from all authors. James Chua contributed to early exploratory experiments, trained the emergently misaligned models, assisted with model diffing experiments, and revised the paper. Clément Dumas and Julian Minder implemented the emergent misalignment model diffing experiments. Kit Fraser-Taliente, Subhash Kantamneni, and Euan Ong implemented all experiments involving the Claude Haiku model. Arnab Sen Sharma and Daniel Wen created the classification datasets. Owain Evans and Samuel Marks jointly supervised the project and revised the paper. All authors contributed to discussions and provided feedback throughout the project.
@@ 11 Acknowledgements
We would like to thank Dan Mossing, Abhay Sheshadri, Oliver Daniels, Jack Lindsey, and Daniel Filan for useful discussions and helpful feedback. We thank Andy Arditi for creating the version of the PersonaQA dataset used in this work. AK did this work as part of the MATS Fellowship. JC and OE were supported by grants from Coefficient Giving. We are grateful to Constellation for hosting.
@@ A.2 Group by Length Batching
As we did not implement sequence packing, we instead use group by length batching to minimize padding tokens and improve training efficiency. We group training examples into mega-batches of size $\text{batch\_size}\times\text{window\_size}$ (16 x 20 = 320 examples with our settings), sort by sequence length in descending order within each mega-batch, then flatten back into the full training order. This ensures that examples within each training batch have similar lengths while maintaining sufficient randomness across mega-batches. We found this approach yielded a speedup of approximately 30% with no noticeable change to final train loss or evaluation performance.
@@ A.3 Infrastructure
We generate all activation vectors on the fly during training by temporarily disabling the LoRA adapter when running the target model forward pass. This avoids the storage costs for full sequences of activations.
@@ A.5 Steering Layer Selection
We investigated whether to apply activation steering after transformer layer 0 or layer 1 of the Activation Oracle. Layer indices are 0-based; "after layer N" refers to the residual stream output of model.model.layers[N] in HuggingFace Transformers (for the Qwen3 architecture), after that layer’s attention and MLP computations. While pan2024latentqateachingllmsdecode found layer 0 optimal, we hypothesized that layer 1 might work better with LoRA fine-tuning, as steering at a position further from the embedding layer may be easier for the LoRA adapter to integrate.
@@
Activation Norm Growth During Training. When using the replacement-based steering method from LatentQA, we observed that activations at placeholder token positions had grown to problematic norms by the time they reached the injection layer:
@@ C.1 Secret-Keeping Evaluation Details
Figure 10: Baseline results on secret-keeping tasks on the direct prompts. Figure reproduced from cywiński2025elicitingsecretknowledgelanguage.
For the secret-keeping evaluations in Section 4, we directly use the baseline results reported by cywiński2025elicitingsecretknowledgelanguage. Figure 10 reproduces their main results figure, showing the performance of various white-box and black-box methods across the three secret-keeping tasks. Their best-performing white-box methods were logit lens-based techniques for Taboo and Secret Side Constraint, and SAE-based techniques for User Gender. Their best-performing black-box method across all three tasks was response continuation with a prefilled prompt. Exact numerical results are available in their code repository at https://github.com/cywinski/eliciting-secret-knowledge/blob/main/METRICS.md.
@@ C.4 Evaluation Input Token Selection
Additionally, we found that focusing the the final 10 tokens (which contain the user end of turn and assistant beginning of turn tokens) could slightly improve performance on the gender evaluation, going from 0.970 to 0.995.
This suggests that providing the full sequence is a reasonable default, though having the flexibility to use single-token or short selection inputs can be valuable for specific tasks. However, in real-world auditing applications without ground-truth labels, it may be difficult to sweep token selection as a hyperparameter. In such cases, providing the full sequence of activations is the simplest option and performs well across most settings.